A silver metal reacts with sulfur to form silver sulfide according to the following reaction: 2Ag(s) + S(s) → Ag2S(s)



.identify the limiting reagent if 50.0 g Ag reacts with 10.0 g S


What is the theoretical yield in g of AgaS produced from the reaction?


What is the amount in g of the excess reactant expected to remain after the reaction?



When the reaction occurred, the amount of Ag-S obtained was 45.0g. What is the percent yield of the reaction?



1
Expert's answer
2021-11-05T01:07:23-0400

2Ag(s) + S(s) → Ag2S(s)


For one mole of sulphur there should be two moles of silver for complete reaction.

Moles of silver = 50/108 = 0.46 moles

Moles of elemental sulphur = 10/32 = 0.3125 moles

For 0.3125 moles of sulphur there must be 0.6250mole of silver.

So silver is limiting Reagent.